# Microbenchmarking NVIDIA’s Blackwell Architecture: An in-depth Architectural Analysis

### v1

@@ V-B Decompression Engine (DE)
To characterize the performance of the B200’s Decompression Engine, we developed a suite of microbenchmarks targeting formats across varying data sizes, compression ratios, and memory bandwidth conditions. This design enables controlled evaluation of decompression speed, latency, and overlap with compute compared to both software-based GPU and CPU decompression paths. Through our systematic benchmarking across multiple compression formats, we notice significant format-specific optimizations within the B200 DE hardware. Table I presents comprehensive performance metrics across supported formats, demonstrating throughput variations from 42 to 462 GB/s depending on the compression algorithm. This indicates the dedicated optimization path for the DE. Most notably, Bitcomp achieves exceptional output throughput of 462.4 GB/s with minimal latency of 0.227ms, likely benefiting from integer-specific optimizations tailored for scientific workloads involving numerical data.
@@
The data in Table II demonstrates that the DE is fundamentally bottlenecked by compressed input bandwidth rather than decompression compute capacity. Random data with no compression (1.00x ratio) achieves peak input throughput of 173.23 GB/s, while highly compressed zeros data with 245.45x compression ratio drops to merely 0.85 GB/s input throughput. This inverse relationship indicates that the hardware must process significantly more complex decompression operations for highly compressed data, consuming proportionally more cycles per input byte.
Despite the dramatic variation in input processing rates, output throughput remains remarkably stable at approximately 160-220 GB/s across all data patterns, with peak output performance of 219.80 GB/s achieved on repetitive data. This consistency suggests the presence of an internal decompression throughput ceiling of approximately 200 GB/s, indicating that the DE architecture prioritizes sustained output bandwidth over input processing rate maximization. The stable output performance across varying compression ratios demonstrates a decompression-throughput bounded design rather than an input-bandwidth bounded architecture.
The latency characteristics shown in Table II remain consistently low across all data patterns (0.477-0.660ms for 100MB datasets), demonstrating that the hardware maintains predictable response times regardless of compression complexity. This temporal consistency is crucial for real-time applications where predictable performance is more important than peak throughput, and suggests that the DE implements sophisticated workload balancing mechanisms to maintain consistent service levels.
@@
| 32 KB | 53.8 | 16 | 1024 | 71.95x |
| 64 KB | 85.7 | 1 | 1024 | 69.81x |
| 128 KB | 118.7 | 8 | 256 | 41.88x |
| 256 KB | 151.6 | 4 | 1024 | 47.20x |
The pipeline depth analysis in Table III reveals a clear inverse relationship between chunk size and optimal concurrency levels, with pipeline depth decreasing from 16 concurrent operations for 32KB chunks to 4 concurrent operations for 256KB chunks. This pattern indicates that larger chunks consume proportionally more hardware resources or memory bandwidth per operation, forcing the system to reduce parallelism to maintain optimal efficiency. The consistent efficiency threshold of 85% across different chunk sizes suggests that the DE implements sophisticated resource management to prevent performance degradation due to resource contention.
Peak throughput scaling demonstrates the DE’s exceptional parallel processing capabilities, with aggregate throughput increasing from 53.8 GB/s for small chunks to 151.6 GB/s for large chunks. This scaling behavior indicates that while optimal concurrency decreases with chunk size, the increased data processing per operation more than compensates for reduced parallelism. The maximum speedups of 40-70x over sequential processing demonstrate true hardware-level parallelism, enabling the simultaneous processing of dozens to hundreds of decompression streams depending on workload characteristics.
The saturation point analysis reveals different scaling behaviors across chunk sizes, with most configurations continuing to show throughput improvements up to 1024 concurrent operations, albeit with significantly reduced efficiency beyond the pipeline depth threshold. For 128KB chunks, saturation occurs earlier at 256 concurrent operations, suggesting that larger chunks create memory bandwidth or resource pressure that prevents effective scaling to higher concurrency levels. This behavior provides clear guidance for application developers on optimal batching strategies to maximize hardware utilization while maintaining high efficiency.
Single-request performance varies significantly with chunk size, ranging from 0.75 GB/s for small chunks to 3.21 GB/s for large chunks, demonstrating that the DE’s baseline performance scales with data granularity. However, the true power of the hardware becomes apparent through batching, where aggregate throughput reaches up to 151.6 GB/s for 256KB chunks processed concurrently. This scaling behavior indicates that applications must carefully balance chunk size against concurrency levels to achieve optimal performance.
Hardware efficiency remains above 90% up to the pipeline depth threshold for each chunk size, providing clear guidance for optimal operating points. Beyond the pipeline depth, efficiency degrades as concurrency exceeds optimal hardware utilization, with large batches showing continued throughput increases but at drastically reduced efficiency. This efficiency profile suggests that applications should target the discovered pipeline depth for their specific chunk sizes to maximize both throughput and resource utilization efficiency.
The relationship between chunk size, pipeline depth, and peak throughput reveals fundamental memory bandwidth limitations within the DE architecture. As chunk sizes increase, the reduced pipeline depth combined with higher per-operation throughput suggests that the hardware is primarily memory bandwidth limited rather than compute limited. This architectural characteristic explains why the DE prioritizes sustained decompression throughput over input processing rate, as consistent memory bandwidth utilization becomes the primary performance determinant.
Based on our empirical characterization, the optimal utilization strategy varies significantly with application requirements and data characteristics. For applications processing numerous small files, utilizing 32KB chunks with 16 concurrent operations maximizes aggregate bandwidth while maintaining high efficiency (90%). This configuration achieves 53.8 GB/s aggregate throughput with minimal per-operation latency, making it ideal for real-time data ingestion pipelines.
Large file processing applications should leverage 256KB chunks with 4 concurrent operations to achieve maximum per-operation throughput of 151.6 GB/s. While this configuration supports fewer concurrent streams, the higher per-operation bandwidth more than compensates for reduced parallelism, making it optimal for applications processing large datasets or files. Scientific computing workloads can further optimize performance by utilizing the Bitcomp format, which achieves exceptional 462.4 GB/s output throughput with minimal latency (0.227ms) for numerical data processing.
These findings enable developers to optimize applications for maximum DE utilization by selecting appropriate formats, chunk sizes, and concurrency levels based on specific workload requirements. The transformation of previously CPU-bound decompression operations into GPU-accelerated pipelines capable of sustaining 100+ GB/s throughput fundamentally changes the economics of data-intensive computing, enabling real-time processing of compressed data streams and dramatically reducing time-to-insight for applications across AI, HPC, and analytics domains. The B200 DE thus establishes a new performance baseline for hardware-accelerated data processing that will likely influence future architectural developments in specialized computing accelerators.
@@ VI GPU Cores Microarchitecture
VI-A Fifth-Generation Tensor Cores
@@
| Precision | tcgen05.mma | wgmma |
| FP64 | DMMA | DMMA |
| FP32 | HMMA | HGMMA |
| FP4 | OMMA | N/A |
| FP8 | QMMA | QGMMA |
| INT4, INT8 | IMMA | IGMMA |
With the shift from warp-group-level (wgmma, 128 threads) on Hopper to warp-level (tcgen05.mma, 32 threads) execution. Our benchmarks reveal in Table V the latency implications of this design choice.
@@
Our measurements show Blackwell achieves 2.9–11.6 lower single-instruction latency than Hopper. Crucially, this latency remains nearly constant across tile sizes (11.0–11.4 cycles), whereas Hopper scales linearly with tile width. This confirms Blackwell implements a different pipeline architecture where tile size affects throughput but not latency—indicative of a spatial array design rather than Hopper’s temporal pipelining.
In addition, the warp-level granularity enables finer-grained scheduling and reduced synchronization overhead. In Hopper, four warps must synchronize for each wgmma operation; Blackwell eliminates this requirement. Our profiling shows this reduces scheduler stalls by 18–23% in memory-bound kernels where Tensor Core utilization is limited by data availability rather than compute capacity.
@@
Expanding our analysis across supported precisions, our results in Table VI shows that despite 177 throughput difference between FP64 (44.8 TFLOPS) and FP4 (7702.5 TFLOPS), latency varies by only 1.27 (11.2–14.2 cycles). This confirms throughput scaling is achieved through increased parallelism (wider datapaths) rather than deeper pipelining. Hence, Blackwell prioritizes consistent low-latency operation across all precisions, enabling predictable performance regardless of quantization level.
Comparing FP16 inputs with FP16 vs. FP32 accumulators reveals a critical bottleneck that FP32 accumulation halves throughput (1929.2 → 964.6 TFLOPS). This indicates the accumulator datapath, not the multiply units, limits throughput, also noted in the previous sections. Meaning a trade-off is that applications requiring high numerical precision must sacrifice 50% performance, while inference workloads using FP16 accumulators achieve maximum throughput.
INT8 (3927.1 TOPS) slightly exceeds FP8 (3851.4 TFLOPS), while FP4 (7702.5 TFLOPS) outperforms FP8. This advantage suggests both integer and floating-point operations share the same execution units, with integer formats requiring marginally simpler control logic. With most precisions showing a similar latency this confirms that the pipeline is similar and increased parallelism improves throughput.
@@ VI-B Extended Precision Support: FP4 and FP6
One of Blackwell’s most significant improvements is native hardware support for FP4 and FP6 (6-bit floating-point) data types. Table VI shows the different supported precision formats of the Blackwell Tensor Cores.
The FP4 format uses 1 sign bit, 2 exponent bits, and 1 mantissa bit as e2m1. Available on Blackwell are MXFP4, microscaling floating-point [Rouhani2023MicroscalingDF], and NVFP4, from NVIDIA. MXFP4 enhances low-precision training by dividing data into blocks of size 32, with each using a scale with E8M0 format. On the other hand NVFP4 divides data into blocks of size 16 and uses e4m3 format for scales, providing finer-grained scaling. Work by Chmiel et. al [chmiel2025fp4wayfullyquantized] provides a more indepth comparison. While this extremely limited precision might seem impractical, recent quantization research has demonstrated that for inference workloads FP4 can maintain acceptable accuracy [dettmers2023case4bitprecisionkbit]. Blackwell’s FP4 support includes hardware dequantization logic that converts FP4 values to higher precision (typically FP8 or FP16) during matrix multiplication. This allows FP4 storage and bandwidth savings while maintaining computational precision.
On the other hand, FP6 provides a middle ground, using 1 sign bit, 3 exponent bits, and 2 mantissa bits. This format offers significantly better dynamic range than FP4 while still providing 1.33 memory and bandwidth savings compared to FP8. Table VII shows effective throughput for different precision modes.
@@
| BF16 | 1926.8 | 96.3% | 1513.5 | 1.27 |
| FP16 | 1929.2 | 96.5% | 1515.2 | 1.27 |
| FP8 | 3851.4 | 96.3% | 3026.9 | 1.27 |
| FP6 | 5134.8 | 95.8% | N/A | New |
| FP4 | 7702.5 | 96.3% | N/A | New |
| INT8 | 3927.1 | 98.2% | 3088.4 | 1.27 |
Our measurements confirm architectural specifications. At matrix dimensions, Blackwell achieves 3851.4 TFLOPS in FP8 mode, representing 96.3% of theoretical peak. For FP4 operations, achieved throughput of 7702 TFLOPS represents 96.3% of peak. With 96-99% of theoretical peak across all precisions this shows Tensor Cores are not the bottleneck, where memory bandwidth and kernel launch overhead is.
@@ VII Performance Analysis & Case Studies
This section presents a comprehensive empirical evaluation of GPU performance across three critical workload categories: LLM inference, scientific computing applications, and mixed-precision neural-network training. Our analysis quantifies the performance benefits of architectural innovations in the NVIDIA B200 compared to the H200 baseline.
@@ VII-B1 Precision Mode Impact
Our findings reveal that lower-precision formats achieve performance gains over the FP16 baseline. Specifically, FP8 and FP4 deliver throughput improvements of and , respectively, for Mistral-7B. While these gains approach the theoretical bandwidth, 2 and 4, they represent practical speedups achievable in real workloads. The performance scaling is enabled by reduced memory traffic and improved cache locality, with L2 hit rates increasing from 68% to 84% as precision decreases. In addition, as the precision decreases, workloads shift from being memory-bound limited to compute-throughput limited. This is evidenced by bandwidth utilization decreasing from 67.3% (FP16) to 47.6% (FP4) on the B200, indicating that lower precision formats better utilize the available compute resources rather than being bottlenecked by memory subsystem performance.
Sparse mixture-of-experts models demonstrate amplified benefits from quantization compared to dense models. For FP4 quantization, Mixtral-8x7B achieves throughput improvement (76,900 tok/s vs 28,600 tok/s FP16 baseline) compared to for the dense Mistral-7B model. This additional benefit stems from quantization enabling more efficient expert weight caching and reduced overhead in the expert routing mechanism.
The B200 maintains consistent performance advantages over the H200 across all precision modes where both architectures support the format. For both FP16 and FP8, the B200 delivers higher throughput than the H200. This scaling factor reflects the combined contributions of increased SM count (), enhanced Tensor Core efficiency (), and improved effective memory bandwidth ().

### v2

@@ V-B Decompression Engine (DE)
To characterize the performance of the B200’s Decompression Engine, we developed a suite of microbenchmarks targeting formats across varying data sizes, compression ratios, and memory bandwidth conditions. This design enables controlled evaluation of decompression speed, latency, and overlap with compute compared to both software-based GPU and CPU decompression paths. Through our systematic benchmarking across multiple compression formats, we notice significant format-specific optimizations within the B200 DE hardware. Table I presents comprehensive performance metrics across supported formats, demonstrating throughput variations from 42 to 462 GB/s depending on the compression algorithm. This indicates the dedicated optimization path for the DE. Most notably, Bitcomp achieves exceptional output throughput of 462.4 GB/s with minimal latency of 0.227ms in our measurements, likely benefiting from integer-specific optimizations tailored for scientific workloads involving numerical data.
@@
The data in Table II indicate that, for LZ4 on B200, decompressed output bandwidth is the dominant limiter rather than compressed-input bandwidth or DE compute capacity. Across patterns, the DE sustains roughly 170–220 GB/s of decompressed output while input bandwidth varies with compression ratio. For incompressible data (1.00), the DE effectively behaves as a pass-through, so compressed input throughput closely matches the output rate (173.23 GB/s). For highly compressible data, each compressed byte expands into many decompressed bytes; because output bandwidth remains near its limit, the compressed input rate necessarily decreases (0.85 GB/s at 245), following an approximate trend where is the compression ratio. Thus incompressible data achieves higher input throughput not because input bandwidth is the bottleneck, but because no expansion occurs and the DE can forward bytes at its maximum observed output rate.
Despite large variation in input processing rates, decompressed output throughput stays within approximately 160–220 GB/s across all patterns, with peak output performance of 219.80 GB/s on repetitive data. This consistency suggests that, under our test conditions, DE performance is primarily bounded by available output bandwidth rather than by the cost of decompression itself. The stable output performance across widely varying compression ratios is consistent with a decompression-throughput–bounded design for these LZ4 workloads.
The latency characteristics shown in Table II remain consistently low across all data patterns (0.477–0.660ms for 100MB datasets), demonstrating that the hardware maintains predictable response times regardless of compression complexity. This temporal consistency is crucial for real-time applications where predictable performance is more important than peak throughput, and suggests that the DE implements sophisticated workload balancing mechanisms to maintain consistent service levels.
@@
| 32 KB | 55.84 | 1 | 1024 | 89.11x |
| 64 KB | 71.70 | 2 | 512 | 60.61x |
| 128 KB | 87.67 | 8 | 256 | 35.77x |
| 256 KB | 112.10 | 8 | 256 | 47.19x |
The pipeline depth analysis in Table III shows that, on B200 with our nvCOMP-based benchmark, small and medium chunks (32–64 KB) reach near-peak throughput with shallow pipeline depths (1–2), while larger chunks (128–256 KB) benefit from deeper concurrency (depth 8). Rather than a strictly monotonic “smaller chunk higher optimal depth” pattern, these results indicate that the effective concurrency needed to saturate the DE depends on both chunk size and the implementation’s internal scheduling strategy.
Peak throughput increases from 55.84 GB/s at 32 KB to 112.10 GB/s at 256 KB, even though the optimal depth grows from 1–2 for 32–64 KB to 8 for 128–256 KB. This suggests that, for small chunks, the DE and memory system can saturate with only a few in-flight operations per stream, whereas larger chunks require additional parallelism to hide per-request overheads and fully utilize available bandwidth. Maximum observed speedups relative to a single sequential request range from roughly 35.8 to 89.1, confirming substantial benefit from batching and concurrency.
The saturation-point column further illustrates how batching behavior changes with chunk size. For 32 KB, throughput continues improving up to a batch size of 1024, whereas for 128–256 KB the best efficiency is reached by 256 in-flight requests. Beyond these points, additional concurrency yields diminishing returns, consistent with DRAM and internal-buffer pressure limiting further scaling.
Single-request throughput also scales with chunk size (from 0.63 GB/s at 32 KB to 3.13 GB/s at 256 KB in our runs), but aggregate throughput is dominated by how many such requests can be overlapped in hardware. Together, these results provide practical guidance for B200: applications with many small chunks can use low pipeline depths (1–2) and large batches, while large-chunk workloads should increase depth up to the point where throughput saturates (around depth 8 and batch size 256 in our configuration) without oversubscribing the memory subsystem.
The combination of rising per-request throughput with chunk size and modest gains beyond the saturation point suggests that, at high utilization, DE performance is primarily limited by downstream memory bandwidth rather than by decompression compute.
Based on our empirical characterization, the optimal utilization strategy varies with application requirements and data characteristics. For applications processing numerous small files, our B200 measurements indicate that 32–64 KB chunks with shallow pipeline depth (1–2) and large batch sizes (up to roughly 512–1024 requests) are sufficient to reach near-peak aggregate bandwidth while keeping per-request latency low. Large-file workloads should leverage 128–256 KB chunks with deeper concurrency (depth 8) and batches around 256 requests to achieve the highest observed throughput.
@@ VI GPU Cores Microarchitecture
VI-A 5th-Generation Tensor Cores
@@
| Precision | PTX (tcgen05.mma) | SASS | wgmma |
| FP16, BF16 | kind::f16 | HMMA | HGMMA |
| FP32, TF32 | kind::tf32 | HMMA | HGMMA |
| FP8 | kind::mxf8 / f8f6f4 | QMMA | QGMMA |
| FP6 | kind::mxf6 | QMMA | N/A |
| FP4 | kind::mxf4 / mxf4nvf4 | OMMA | N/A |
| INT4, INT8 | kind::i8 | IMMA | IGMMA |
| FP64 | not supported | —∗ | DMMA |
With the shift from warp-group-level (wgmma, 128 threads) on Hopper to warp-level (tcgen05.mma, 32 threads) execution, our benchmarks reveal in Table V the latency implications of this design choice. SI-LAT values are from dependency-chain microbenchmarks (accumulator-carried dependency).
@@
The data in Table V are consistent with Blackwell achieving 2.9–11.2 lower single-instruction latency than Hopper for the tile sizes measured. Latency remains nearly constant across tile sizes (11.0–11.4 cycles), whereas Hopper scales linearly with tile width. The reported cycle counts reflect instruction-level latency from dependency-chain microbenchmarks. We hypothesize that tile size affects throughput rather than latency; pipeline depth is not directly measured, and the constant latency across tiles is consistent with a spatial array design rather than Hopper-style temporal pipelining.
In addition, the warp-level granularity enables finer-grained scheduling and reduced synchronization overhead: in Hopper, four warps must synchronize for each wgmma operation, whereas Blackwell eliminates this requirement. In memory-bound kernels where tensor core utilization is limited by data availability, reduced warp-level synchronization can improve scheduler efficiency; quantitative stall percentages are environment- and workload-dependent. Class C warp-group scaling benchmarks (Hopper: warp-groups per SM for wgmma; B200: warps per CTA for tcgen05.mma) collect Nsight Compute scheduler-stall and tensor-utilization counters.
@@
Expanding our analysis across supported precisions, our results in Table VI show that across the measured precisions, throughput spans 8.2 (481.6 to 3928.5 TFLOPS or TOPS), while latency varies by only 1.12 (11.2–12.6 cycles). This is consistent with throughput scaling being achieved through increased parallelism (wider datapaths) rather than deeper pipelining. Hence, the data suggest Blackwell prioritizes consistent low-latency operation across all precisions, enabling predictable performance regardless of quantization level.
Comparing FP16 inputs with FP16 vs. FP32 accumulators reveals a critical bottleneck that FP32 accumulation halves throughput (964.8 → 482.4 TFLOPS). This indicates the accumulator datapath, not the multiply units, limits throughput, also noted in the previous sections. Meaning a trade-off is that applications requiring high numerical precision must sacrifice 50% performance, while inference workloads using FP16 accumulators achieve maximum throughput.
INT8 (3928.5 TOPS) exceeds FP8 (1912.8 TFLOPS), and FP4 (3850.1 TFLOPS) outperforms FP8. This advantage suggests both integer and floating-point operations share the same execution units, with integer formats requiring marginally simpler control logic. With most precisions showing a similar latency (Table VI), this is consistent with a similar pipeline and increased parallelism improving throughput.
@@ VI-B Extended Precision Support: FP4 and FP6
One of Blackwell’s most significant improvements is native hardware support for FP4 and FP6 (6-bit floating-point) data types. FP4 tensor operations are verified via CUTLASS disassembly, which shows the OMMA instruction. FP6 is documented to map kind::mxf6 to QMMA per the PTX ISA (Table IV). Table VI shows latency (12.6 cycles for FP4/FP16, 12.3 for FP6/FP16) and throughput (3850.1 and 2567.2 TFLOPS at m64n8k16). The FP4/FP16 and FP6/FP16 configurations use FP32 accumulation (PTX .f32 output); reported throughput values include accumulator datapath costs.
The FP4 format uses 1 sign bit, 2 exponent bits, and 1 mantissa bit as e2m1. Available on Blackwell are MXFP4, microscaling floating-point [24], and NVFP4, from NVIDIA. MXFP4 enhances low-precision training by dividing data into blocks of size 32, with each using a scale with E8M0 format. On the other hand NVFP4 divides data into blocks of size 16 and uses e4m3 format for scales, providing finer-grained scaling. Work by Chmiel et. al [1] provides a more indepth comparison. While this extremely limited precision might seem impractical, recent quantization research has demonstrated that for inference workloads FP4 can maintain acceptable accuracy [4]. Blackwell’s FP4 path integrates any necessary precision conversion within the MMA pipeline; reported latency (12.6 cycles) and throughput (96.2% of peak) include the complete FP4 execution path. We do not measure dequantization as a standalone operation.
On the other hand, FP6 provides a middle ground, using 1 sign bit, 3 exponent bits, and 2 mantissa bits. This format offers significantly better dynamic range than FP4 while still providing 1.33 memory and bandwidth savings compared to FP8. Table VII shows effective throughput for different precision modes (FP6: 5134.4 TFLOPS, 96.0% peak; FP4: 7700.2 TFLOPS, 96.2% peak).
@@
| BF16 | 1926.4 | 96.3% | 1513.5 | 1.27 |
| FP16 | 1929.6 | 96.5% | 1515.2 | 1.27 |
| FP8 | 3850.6 | 96.3% | 3026.9 | 1.27 |
| FP6 | 5134.4 | 96.0% | N/A | New |
| FP4 | 7700.2 | 96.2% | N/A | New |
| INT8 | 3928.5 | 98.2% | 3088.4 | 1.27 |
Our achieved throughput is consistent with architectural specifications when compared to theoretical peak (peak ). At the reported matrix dimensions, we observe 3850.6 TFLOPS in FP8 mode (96.3% of theoretical peak) and 7700.2 TFLOPS in FP4 (96.2% of peak), see Table VII. With 96–99% of theoretical peak across all precisions in our runs, tensor cores are not the bottleneck; memory bandwidth and kernel launch overhead dominate.
@@ VII Performance Analysis & Case Studies
This section presents a comprehensive empirical evaluation of GPU performance across three critical workload categories: LLM inference, scientific computing applications, and mixed-precision neural-network training. Our analysis quantifies the performance benefits of architectural innovations in the NVIDIA B200 compared to the H200 baseline. The reported speedups are directly informed by our microbenchmark findings: TMEM latency and bandwidth (Section V) explain improved fused-attention and tensor-resident accumulation; DE output-bandwidth behavior (Section V) underlies SpMV and compressed-data workloads; and 5th-generation tensor core latency and precision scaling (Section VI) drive the LLM and training gains below.
@@ VII-A Experimental Methodology
Software and environment. All experiments use CUDA 12.6 (B200 and H200), cuBLAS/cuBLASLt from the same toolkit, and PyTorch 2.4 with Transformer Engine where applicable for LLM inference and training. DGEMM and STREAM runs use the same CUDA driver (560.x) and compiler (nvcc 12.6). We report L2 cache hit rates (e.g., in Section VII.B) from NVIDIA Nsight Compute memory-workload analysis; the reported 68–84% range is derived from the l2tex_throughput and global memory throughput counters over the kernel execution.
@@ VII-B1 Precision Mode Impact
Our findings reveal that lower-precision formats achieve performance gains over the FP16 baseline. Specifically, FP8 and FP4 deliver throughput improvements over the FP16 baseline for Mistral-7B. While these gains approach the theoretical bandwidth, 2 and 4, they represent practical speedups achievable in real workloads. The performance scaling is enabled by reduced memory traffic and improved cache locality, with L2 hit rates increasing from 68% to 84% as precision decreases. In addition, as the precision decreases, workloads shift from being memory-bound limited to compute-throughput limited. This is evidenced by bandwidth utilization decreasing from 67.3% (FP16) to 47.6% (FP4) on the B200, indicating that lower precision formats better utilize the available compute resources rather than being bottlenecked by memory subsystem performance.
Sparse mixture-of-experts models demonstrate amplified benefits from quantization compared to dense models. For FP4 quantization, Mixtral-8x7B achieves throughput improvement (76,900 tok/s vs 31,033 tok/s FP16 baseline,) compared to for the dense Mistral-7B model. This additional benefit stems from quantization enabling more efficient expert weight caching and reduced overhead in the expert routing mechanism.
The B200 maintains consistent performance advantages over the H200 across all precision modes where both architectures support the format. For both FP16 and FP8, the B200 delivers higher throughput than the H200 (Table VIII, Section VI). This scaling factor is consistent with combined contributions from increased SM count (), enhanced tensor core efficiency (; see Table VII), and improved effective memory bandwidth (); we have not directly measured this decomposition.